The time a basketball player spends in the air when shooting a basket is called "the hang time." The vertical leap L measured in feet is related to the hang time "t" measured in sefconds by the equation L=4t^2. A typical athlete has a vertical leap of 1 1/2 to 2 feet; the best male jumpers attain heights of 3 1/2 to 4 feet. What is the hang time of an athlete who has a vertical leap of 2.8 feet?
Answer by smik(40)

We know that L = 4t^2, and we want to find 't' for an athlete who has a value of 2.8 for L. SO:


2.8 = 4t^2 (divide both sides by 4)
0.7 = t^2 (take the square root of both sides)
t = 0.7^(1/2) = 0.837


An athlete who has a vertical leap of 2.8 feet would have a hang time of approximately 0.837 seconds.